A bag of marbles contains 6 blue, 7 red, and 8 orange marbles. What is the probability of picking a red marble and then an orange marble if you do not replace the first marble?
Answer by jorel555(1290)

The probability of picking a red marble, then an orange one, without replacement, is:
7/21 x 8/20=56/420=.133333333 . ☺☺☺☺